OpenStudy (anonymous):

A fair coin tossed 8 times. Find the probability of getting exactly 3 Heads

hartnn (hartnn):

how much is the probability of getting one head ?

OpenStudy (anonymous):

it doesnt say

hartnn (hartnn):

let that probability be p, exactly 3 heads means , exactly 3 heads and 5 tails. so the probability of getting exactly 3 Heads will be p^3 (1-p)^5 = ?

hartnn (hartnn):

the probability of getting one head in one coin toss is ?

hartnn (hartnn):

it won't say, u need to know it

OpenStudy (anonymous):

it says to use a table given but its confusing to read

hartnn (hartnn):

which table ?

OpenStudy (anonymous):

its in my book its a binomial table

hartnn (hartnn):

does the question say u need to use that or you can do it without the table also ?

OpenStudy (anonymous):

it says i can use it for binomial distribution

hartnn (hartnn):

u still have to figure :the probability of getting one head in one coin toss is ?

hartnn (hartnn):

one coin toss can either produce a head or tail. so there is equal probability of getting those which is ?

OpenStudy (anonymous):

.5?

hartnn (hartnn):

that is correct! so p= 0.5 using binomial distribution u get your probability as p^3 (1-p)^5 = (0.5)^3 (1-0.5)^5 = ?

OpenStudy (anonymous):

what symbol is ^?

hartnn (hartnn):

\(\huge (0.5)^3 (1-0.5)^5=?\)

hartnn (hartnn):

or if u have to use table: you can go to a Binomial Distribution Table and find the solution, values for n=8 and p=0.5, read from the section the probability of 3 which is ?

OpenStudy (anonymous):

.12?

hartnn (hartnn):

u have options/choices ?

OpenStudy (anonymous):

well it doesnt say i have to use the table

OpenStudy (anonymous):

would my answer be correct?

hartnn (hartnn):

i don't have the table, just look up for values for n=8 and p=0.5, read from the section the probability of 3, that would be your answer

OpenStudy (anonymous):

ok thanks, appreciate the help

hartnn (hartnn):

welcome :)

OpenStudy (raden):

but, the sum of ways we need look also, right @hartnn ? so, the sum of way be equal 8!/(3!*5!) ?

hartnn (hartnn):

would that be like multiplying 8C5 , yes, i thought it before....but since cst had a table, i then resorted to using table only....

hartnn (hartnn):

8C5 *p^3 (1-p)^5 will give the same value mentioned in table

hartnn (hartnn):

which would be like 0.21875, if i calculated it properly....

OpenStudy (raden):

yea, like that i mean